OpenStudy (anonymous):

Goodness, what about this one f(x)=cos(arcsin(9x))

OpenStudy (anonymous):

arcsin(9x) is easy, but i'm not sure what to do with the cos out front

OpenStudy (anonymous):

you mean write it as a function of \(x\) with no trig?

OpenStudy (anonymous):

Yeah, find the derivative

OpenStudy (anonymous):

start with \[\cos(\arcsin(9x))=\frac{1}{\sqrt{1+81x^2}}\] and take the derivative of that

OpenStudy (anonymous):

Ummmm, \[\frac{ 1 }{ \sqrt{162x} }\]?

OpenStudy (anonymous):

or do i use the quotient rule?

zepdrix (zepdrix):

\[\Large\rm \arcsin(9x)=\theta\qquad\to\qquad \sin \theta=9x\]

zepdrix (zepdrix):

Pythagorean Theorem to find this missing side.